OpenStudy (anonymous):

A computer base unit of mass 7.5 kg is dragged along a smooth desk. If the normal contact force is 23N and the tension in the arm of the person dragging it acts at 23 degrees to the horizontal, then what is the total tension in the person's arms?

OpenStudy (anonymous):

 The arrow shows the direction of tension acting on the body. The arrow above is the normal force, the arrow below is gravitation. Since the body is in vertical equilibrium [It's not accelerating up or down], we can say that \[T \sin (\theta) + N = mg\] where m is the mass of the body, N is the normal contact force. We know N, theta, m and g. So we can find T.

OpenStudy (anonymous):

The answer comes to about 129.5 Newtons.

OpenStudy (anonymous):

ts\[tsin \theta+n=mg where n=contact force m =mass comp base\] calculate..